OpenStudy (anonymous):

Find the Difference. 3x/3x^2+10x-8 (minus) 9x/5x^2+11x-36

OpenStudy (anonymous):

I really need the answer

OpenStudy (callisto):

\[\frac{3x}{3x^2+10x-8} - \frac{9x}{5x^2+11x-36}\]\[=\frac{3x}{(3x-2)(x+4)} - \frac{9x}{(5x-9)(x+4)}\]\[=\frac{3x(5x-9)}{(3x-2)(x+4)(5x-9)} - \frac{9x(3x-2)}{(5x-9)(x+4)(3x-2)}\]\[=\frac{3x(5x-9)-9x(3x-2)}{(5x-9)(x+4)(3x-2)}\]\[=\frac{15x^2-27x-27x^2+18x}{(5x-9)(x+4)(3x-2)}\]\[=\frac{-12x^2 - 9x}{(5x-9)(x+4)(3x-2)}\]\[=\frac{-3x(4x + 3)}{(5x-9)(x+4)(3x-2)}\]

OpenStudy (anonymous):

Is that the answer? Thank you.

OpenStudy (callisto):

It can be, or may not be. You can further simplify it..